What is the magnitude of the angular momentum of a 900 g, 2.0 m long bar rotating at 200 rpm about an axis through its center? Assume the bar is thin.
Answer #1

angular momentum L = I*ω

m = 900 g = 0.900 kg

l = 2 m

ω = 200 rpm = 200*2pi/60 =20.943951 rad/s

I = m*l2 /12 = 0.900*2^2/12 =0.3 kg.m2

SO

angular momentum L = I*ω = 0.3*20.943951 = 6.2831853 kg-m2/sec

*************************************************************************
Goodluck Comment in case any doubt, will reply for sure..